OpenStudy (anonymous):

The equations x=1+2cos(theta) y=-2+3sin(theta) parameterize what curve in the xy-plane? Determine its Cartesian equation.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Ok well do you have any ideas about how to think about this? What would the curve x = cos(theta) y = sin(theta) parameterize?

OpenStudy (anonymous):

I'm really not sure, I'm getting tripped up by "parameterize." I know cos curve and sin curve are the same but just start on different intervals (cos starts at y=1 sin start at x=0 ?)

OpenStudy (anonymous):

Ah. Ok

OpenStudy (anonymous):

Well lets imagine that we have something simple like y = sin(theta) x = cos(theta) Then when we plug in theta = 0 we get x = 1, y = 0 or the point (1,0) If we plug in \(\pi \over 6\) we get x = \(\sqrt{3} \over 2\), y = \(1 \over 2\) If we plug in \(\pi \over 4\) we get x = \(\sqrt{2} \over 2\), y = \(\sqrt{2} \over 2\) If we plug in \(\pi \over 3\) we get x = \(1 \over 2\), y = \(\sqrt{3} \over 2\) If we plug in \(\pi \over 2\) we get x = 0, y = 1 etc.

OpenStudy (anonymous):

Do these numbers look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we would take the values we plugged in to eqaute the value for the curve? (sorry I'm very lost thus far in Calc III) The numbers look familiar from using them in calc II with trig sub

OpenStudy (anonymous):

x = cos(theta), y = sin(theta) is the standard parameterization of a circle. As we plug in various values for theta we will get corrisponding x and y points that describe a circle around the origin with radius 1.

OpenStudy (anonymous):

If you imagine a radius fixed at the origin making an angle theta from the x axis the x and y value of the outer tip will be the x and y values we get from our parameterization.

OpenStudy (anonymous):

oh, I think I'm understanding it a little better

OpenStudy (anonymous):

So now imagine that we had instead x = 2 + cos(theta) y = 3 + sin(theta) What would this change?

OpenStudy (anonymous):

the different values would just be shifted by the coefficient right?

OpenStudy (anonymous):

Would we still be describing a circle?

OpenStudy (anonymous):

Would it still have radius 1?

OpenStudy (anonymous):

Would it still be centered at the origin?

OpenStudy (anonymous):

no it would be greater

OpenStudy (anonymous):

ohhh and the origin would be shifted as well??

OpenStudy (anonymous):

Well try plugging in a few common points \(\pi, 0, \pi/2, \pi/4, etc\). What do you get for some of the x/y values?

OpenStudy (anonymous):

1 + 2cos(0) = 3 -2 + 3sin(0) = -2 1+2cos(pi/2) =1 -2 + 3 sin(pi/2) = -1

OpenStudy (anonymous):

I meant plug into my new version of the function.

OpenStudy (anonymous):

Just to see how changing different pieces can change the function

OpenStudy (anonymous):

oh sorry, hold on

OpenStudy (anonymous):

they seem to move the same only deviate by the number being added?

OpenStudy (anonymous):

cos (pi/4) and sin (pi/4) are equivalent?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So what does that tell you about adding a constant value to the original function? What did it do?

OpenStudy (anonymous):

so only the radius is changing? maybe :)

OpenStudy (anonymous):

Did the radius change? x = 2 + cos(theta) y = 3 + sin(theta) theta x,y At 0 3,3 At \(\pi \) 1,3 At \( \pi/2\) 2,4 At \(3\pi/2\) 2,2

OpenStudy (anonymous):

no.. is it the length of the curve thats changing while the radius remains constant? thats just a wild guess :(

OpenStudy (anonymous):

No.. The radius is 1. All that has happened is we've changed where we're drawing the circle. Now instead of being centered around 0,0 we are centered around 2,3. It's still just an ordinary circle.

OpenStudy (anonymous):

oh!!! sorry

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so its a constant circle as long as cos(theta) and sin(theta) are unchanged and anything added it too that shifts the origin?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

Now lets watch what happens when we put a coefficient on the cos and sin

OpenStudy (anonymous):

genius!

OpenStudy (anonymous):

x = 2 + 3cos(theta) y = 3 + 3sin(theta)

OpenStudy (anonymous):

makes the radius bigger

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

And finally, if we make the sin and cos have different coefficients. we will end up having one grow faster than the other which will skew our circle making it an ellipse.

OpenStudy (anonymous):

x=1+2cos(theta) y=-2+3sin(theta)

OpenStudy (anonymous):

ahh, its starting to make sense!

OpenStudy (anonymous):

The length of the major axis being 6 (in the y direction) and the minor axis being 4 (in the x direction) which are twice the respective radial coefficients.

OpenStudy (anonymous):

and still centered at 1,-2

OpenStudy (anonymous):

okay, so when it asks to "parameterize" it would be x=1 and y = -2 ?

OpenStudy (anonymous):

No. It gave the parameterization. It's asking for the Cartesian equation.

OpenStudy (anonymous):

so would Cartesian equation be the eq of a circle (x - a)^2 + (y - b)^2 = r^2

OpenStudy (anonymous):

using x=1 y = -2

OpenStudy (anonymous):

No, it'd be the equation of an ellipse.. \[{(x-h)^2 \over a^2} + {(y-k)^2 \over b^2} = 1\] Where h,k is the center, and a and b are half the lengths of the major and minor axis of the ellipse. In your case h = 1, k = -2, a = 2, b = 3

OpenStudy (anonymous):

Yes, because you just said it was an elipse too, blahh. I wish this stuff would click easier

OpenStudy (anonymous):

you are a genius though! I appreciate the help

OpenStudy (anonymous):

well this is stuff you should have covered in pre-calc.

OpenStudy (anonymous):

so just got to refresh =)

OpenStudy (anonymous):

i never took precalc ! I went straight into calc I. I should def find a resource and review it.

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

There are about 7 videos that are pretty short but should be a good intro.

OpenStudy (anonymous):

excellent! I've watched some of their videos in the past, I will watch these right now

OpenStudy (anonymous):

you went above beyond, many thanks!

OpenStudy (anonymous):

oh, and click the 'good answer' button so it knows you're done with this question =)

OpenStudy (anonymous):

will do!